OpenStudy (anonymous):

CupCaking, a cupcake bakery, uses the equation below to determine the price of cupcakes, where y represents the total price in dollars and x represents the number of cupcakes purchased. Cake Me, another cupcake bakery, uses the table below to determine prices. Number of Cupcakes (x) Price (y) 2 $3.50 3 $5.25 4 $7.00 6 $10.50 Katie bought one cupcake from CupCaking and Sarah bought one cupcake from Cake Me. Which of the following are true?

OpenStudy (anonymous):

A. Sarah is paying the smaller price of $1.25 per cupcake. B. Katie is paying the smaller price of $1.50 per cupcake. C. Katie is paying $0.25 per cupcake more than Sarah. D. They are paying the same price per cupcake. E. Sarah is paying $0.25 per cupcake more than Katie.

OpenStudy (anonymous):

Is This Algebra 1?

Directrix (directrix):

The equation for CupCaking is y = (3/2)*x

Directrix (directrix):

For 1 cupcake, Katie paid y = (3/2)*1 = $1.50 Looking at the Cake Me table of values, observe that when the number of cupcakes increases by 1, the cost increases by $1.75. You may want to verify that. For 1 cupcake, Sarah paid $1.75. With that information, select the two correct options. Post what you get if you'd like for someone to check your answer. @sarahvega